OpenStudy (yisselluna):

Solve the following system of equations by the substitution method. 10x + 10y = 1 x = y - 3 What is the value of y? -29/20 31/20 -3/2

Awolflover1 (awolflover1):

alright lets go step by step ;D

Awolflover1 (awolflover1):

\[x=y-3; 10x+10y=1\]

OpenStudy (yisselluna):

ok

Awolflover1 (awolflover1):

\[10x+10y=1\] \[10(y-3)+10y=1\]

Awolflover1 (awolflover1):

\[10\times10=?\]

OpenStudy (heschamsaleh):

10x + 10y =1 x - y = -3 (multiply by 10 ) then 10x+10y =1 10x-10y =-30 adding both new equations 20x=-29 x=-29/20 then subistute y = 31/20

Awolflover1 (awolflover1):

please go step by step with the user so they won't be confused

OpenStudy (heschamsaleh):

it is my first hour here sorry

OpenStudy (yisselluna):

okk